OpenStudy (anonymous):

Which of the following equations is in slope-intercept form? A. 2y = 3x – 1 B. x = 3y – 1 C. y = –3x D. 9x = 2

OpenStudy (anonymous):

@Avenged7x

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

Answer A

OpenStudy (solomonzelman):

nope

OpenStudy (solomonzelman):

`y=mx+b` Which equation is in the same form ?

OpenStudy (anonymous):

Explain to me @SolomonZelman

OpenStudy (anonymous):

none of them

OpenStudy (anonymous):

Would it be D then

OpenStudy (solomonzelman):

y=mx+b where m and b is a number. m is the slope, and b is the y-intercept of the line.

OpenStudy (anonymous):

c?

OpenStudy (solomonzelman):

No not D, stop guessing !

OpenStudy (anonymous):

Okay, Okay

OpenStudy (solomonzelman):

C is right, but don't guess next time. y=-3x y=mx+b b=0, m=-1

OpenStudy (anonymous):

i wasnt guessing i was looking at y=mx+ b and thought it looked like it

OpenStudy (anonymous):

@SolomonZelman can help me on my Question

OpenStudy (solomonzelman):

Okya... but remember that y-intercept is a form, when you set the equation equal to y.

OpenStudy (solomonzelman):

*is when ...

OpenStudy (anonymous):

Look in your notification

OpenStudy (anonymous):

@SolomonZelman i have 13 minutes left on my tes and i cant fuger the las 3 out will u help

OpenStudy (solomonzelman):

My notifications are off. Half of them are there half of them aren't. IF you are talking to me.

OpenStudy (anonymous):

*test, *Figure

OpenStudy (anonymous):

Wait, why would M=-1? wouldn't M=-3?

OpenStudy (solomonzelman):

Wait a test ? Last time I helped with test questions and mod noticed it I was suspended. It was on my (about) 40th day

OpenStudy (solomonzelman):

And a mod is here... I better gather my stuff and go ;)

OpenStudy (anonymous):

well its actually a pre test

OpenStudy (anonymous):

Equations of Lines Pre-Test Active

OpenStudy (solomonzelman):

okay then... please try to use 1 question per thread. I'll help you with this coming question here, but next time please...

OpenStudy (anonymous):

Which of the following equations is in slope-intercept form? A. 2y = 3x – 1 B. x = 3y – 1 C. y = –3x D. 9x = 2

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

y = –2x + 4 A. x + 2y = 4 B. 2x + y = 4 C. 4x + y = 2 D. 4x + 2y = 0

OpenStudy (solomonzelman):

you asked this just now

OpenStudy (solomonzelman):

Was this one of the times when you click something without knowing what, and it copy pastes your last reply ?

OpenStudy (solomonzelman):

I have that all the time, very annoying.

OpenStudy (anonymous):

yeA

OpenStudy (anonymous):

i have 7 minutes

OpenStudy (solomonzelman):

y = –2x + 4 x + 2y = 4 2x + y = 4 4x + y = 2 4x + 2y = 0 COMPARE to Y=MX+B`

OpenStudy (solomonzelman):

which one do YOU think it is ?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

?????

OpenStudy (solomonzelman):

No

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

5 minutes

OpenStudy (solomonzelman):

\(\Huge\color{black}{ y=\color{red}{m}x+\color{blue}{b}}\) \(\normalsize\color{red}{ \rm slope}\) \(\normalsize\color{blue}{ \rm y-intercept }\)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i because there is only 1 y

OpenStudy (solomonzelman):

There is only one option that is set to `y=... `

OpenStudy (anonymous):

no thats the equeathon that you have to Rewrite the slope-intercept form equation into standard form.of

OpenStudy (anonymous):

y = –2x + 4

OpenStudy (anonymous):

please just give me the answer and then explain it i have 1 min and 40 seconds

OpenStudy (anonymous):

Which formula would you use when you are looking for the slope of a line and given two points on the line? A. point-slope formula B. slope formula C. slope formula and point-slope formula D. slope-intercept formula

OpenStudy (anonymous):

40 seconds

OpenStudy (anonymous):

i failed

OpenStudy (solomonzelman):

I can't give you answers, I don't want to be banned.

OpenStudy (solomonzelman):

I am sorry

OpenStudy (anonymous):

btw i know you cant tell over the internet but im actually not upset at u i know its not u r fault and i do apreciate your assistance

OpenStudy (anonymous):

you do not need to say sorry

OpenStudy (solomonzelman):

I would try to get a tutor to get a better grip on this math before it is too late. Thank you for your cooperation.... have fun !